If two coils are connected in series and have resistance of 400 ohms and 200 ohms with temperature coeffetients of 0.1% and 0.4% respectively at 20 degrees Celsius. What is the temperature coefficient of the combination, and what is the resistance of the combination at 50 degrees Celsius?


1
Expert's answer
2021-05-21T10:42:42-0400

Resistance, "R_1=400\space\Omega"

"\alpha_1=0.001"

Resistance, "R_2=200\space\Omega"

"\alpha_2=0.004"

Equivalent Resistance, "R_{eq}=R_1+R_2=600\space \Omega"

Initial temperature, "T_1=20\degree C"

Final temperature, "T_2=50\degree C"

"\Delta T=30\degree C"

temperature coefficient of the combination,

"\alpha=\dfrac{R_1\alpha_1+R_2\alpha_2}{R_1+R_2}=0.002"


resistance of the combination at 50 degrees Celsius,

"R_f=R_{eq}(1+\alpha\Delta T)"

"R_f=646\space\Omega"